You can travel28 miles in 45 minutes. You want to know the number of minutes you will need to go 59 miles if you travel at the same rate?
Answer by checkley75(3666)

128/45=59/X CROSS MULTIPLY
128X=45*59
128X=2655
X=2655/128
X=20.742 MINUTES TO GO 59 MILES